# GDSG: Graph Diffusion-based Solution Generation for Optimization Problems in MEC Networks

### v1

@@ II-B4 Cost of offloading execution
$\mathcal{C}^{o}_{i}=\alpha_{i}\frac{V_{i}}{F}+(1-\alpha_{i})P_{e}\frac{V_{i}}{F},\ \forall i\in\{1,...,L\}.$ (6)
Therefore, the total offloading cost of $E_{i}$ can be calculated as $D_{i}(\mathcal{C}^{t}_{i}+\frac{\mathcal{C}^{o}_{i}}{A_{i}})$.
II-B5 Additional parameters
Since the computing task has a maximum tolerable delay $\tau_{i}$ [seconds], let $\rho_{i}\in[0,1]$ represent the minimum computational resource allocation ratio required for offloading execution without timeout, and let $\psi_{i}\in\{0,1\}$ represent whether the local execution will timeout or not. Only if the offloading was able to process without a timeout, $\rho_{i}>0$. As a result, $\rho_{i}=0$ if $\frac{I_{i}}{r^{u}_{i}}+\frac{V_{i}}{F}>\tau_{i}$ and $\rho_{i}=\frac{V_{i}}{(\tau_{i}-I_{i}/r^{u}_{i})F}$ if $\frac{I_{i}}{r^{u}_{i}}+\frac{V_{i}}{F}\leq\tau_{i}$. Furthermore, $\psi_{i}=0$ if the local execution will timeout, $\psi_{i}=1$ otherwise. Therefore, we use $(\mathcal{C}^{l}_{i},\mathcal{C}^{t}_{i},\mathcal{C}^{o}_{i},\rho_{i},\psi_{i})$ as the 5-dimensional edge feature in $\mathcal{G}$ to represent computing tasks, communication parameters, and resource settings.
@@ II-C Weighted Cost Minimization Problem
According to Eqs. (4), (5) and (6), the minimization problem of one round offloading optimization can be formulated as
$\mathbb{P}:\underset{\{\mathcal{D},\mathcal{A}\}}{{\rm min}}\ \sum^{L}_{i=1}(1-D_{i})\mathcal{C}^{l}_{i}+D_{i}(\mathcal{C}^{t}_{i}+\frac{\mathcal{C}^{o}_{i}}{A_{i}}),$
@@
Let $\{\mathcal{D}^{*},\mathcal{A}^{*}\}$ be the optimal solution for the given input parameters, the problem $\mathbb{P}$ can be transformed from the original Mix-Integer Non-Linear Programming (MINLP) to a multi-task graph optimization. It is important to note that although $\mathbb{P}$ is formally an MINLP problem, existing solvers are not equipped to handle such a challenging graph-structured problem. For example, GEKKO [48] only accepts vector-based problem and constraint formulations, and its solution time for MINLP problems is excessively long. With the node features that represent node types (e.g., 1 represents a server node and 0 represents a user node) and edge features that explicitly denote objective and constraint, the graph-based problem formulation approach can provide rich information and is transferable to alternative problems.
@@ III Problem Analysis
In this section, we will provide an introduction to the principles and concepts of solution distribution learning, explicitly highlighting the robustness of this learning objective with respect to solution quality, thereby supporting the learning from suboptimal solution data.
@@ III-A Why and How to Learn the Solution Distribution?
For the optimization problem in Eq. (7), the existing methods generally obtain truth input-solution paired samples like $(x,y^{*}),x\in\mathcal{X},y^{*}\in\mathcal{Y}_{x}$, $\mathcal{X}$ and $\mathcal{Y}_{x}$ represent the input parameter space and the feasible solution space for a given $x$, respectively. By referring to and extending the definition of parametrized solution space in DIMES [30], we parametrize the solution space with a continuous vector $\theta\in\mathbb{R}^{N}$, where $N$ represents the dimension of the optimization variable, and estimate the probability of each feasible solution $y$ as
$p_{\theta}(y|x)\propto{\rm exp}(\sum^{N}_{i=1}y_{i}\theta_{i})\ \ s.t.\ y\in\mathcal{Y}_{x},\ x\in\mathcal{X}.$ (8)
@@
where $p_{\theta}$ is an energy function indicating the probability of each solution over the feasible solution space. We first consider the discrete solution space that $y$ is an $N$-dimensional vector with $y_{i}\in\{0,1\}$, and the higher value of $\theta_{i}$ denotes a higher probability that $y_{i}$ is consistent with $y^{*}$. There exists $\theta^{*}$ be the optimal parametrization of the solution space for a given $x\in\mathcal{X}$ if $\theta^{*}$ satisfies $p_{\theta^{*}}(y^{*}|x)=1$.
@@
Accordingly, $\theta$ describes a solution distribution. Taking the discrete binary classification problem as an example, for the optimal solution $y^{*}$, it is parametrized as a binomial distribution, i.e., $\theta^{*}_{i0}=0,\theta^{*}_{i1}=1$ if $y^{*}_{i}=1$, and $\theta^{*}_{i0}=1,\theta^{*}_{i1}=0$ if $y^{*}_{i}=0$, where $\theta^{*}_{i0}$ represents the probability that $y_{i}=0$ and vice versa. Therefore, $(x,y^{*})$ can be replaced by $(x,\theta^{*})$ for learning.
@@
Although learning $\theta^{*}$ can guarantee the results, it is obvious that $\theta^{*}$ is too sharp to be different from discriminative training, which makes parametrization redundant. For the NP-hard problems considered in this paper, it is impossible to efficiently obtain a large number of truth samples for discriminative supervised training. However, suboptimal solutions are much easier to obtain.
@@
Let $(x,y^{\prime})$ be a suboptimal solution for a given $x$, and $(x,\theta^{\prime})$ be the parametrization. To avoid verbosity, we will not introduce the definition of the gap between suboptimal solutions and the optimal solution here and leave it in Sec. III-B. We know that suboptimal parameterizations are not unique. For example, taking $\theta^{\prime}_{1}=(0.9,0.1),\theta^{\prime}_{2}=(0.95,0.05)$ for $\theta^{*}=(1,0)$, where repeated sampling of $p_{\theta^{\prime}_{1}}(y|x)$ or $p_{\theta^{\prime}_{2}}(y|x)$ can easily hit the optimal solution. Therefore, let $Y^{\prime}_{x}$ be the set of suboptimal solutions for $x$, and $\Theta_{x}$ be the set of suboptimal parameterizations.
@@
For the continuous solution space, it is impossible to make the calculated solution completely equal to the $y^{*}$ because of the infinite precision. As a discussion example, we parametrize $y^{*}$ using sharp Laplace distribution. For example, for $y^{*}_{i}$, $\theta^{*}_{i}=y^{*}_{i}$ for ${\rm Laplace}(\theta^{*}_{i},{\rm scale\_factor})$, smaller ${\rm scale\_factor}$ means sharper probability density function and the probability of $y^{*}$ is the largest and tends to 1. Thus, we can obtain the continuous solution space’s $Y^{\prime}_{x}$ and $\Theta_{x}$.
@@
Consequently, when given only a dataset of suboptimal solutions, learning the parametrization of the solution space to sample feasible solutions can lead to the optimal solution. Specifically, the model aims to learn $p(\Theta_{x})$ to obtain $\theta^{\prime}\sim p(\Theta_{x})$ and sample feasible solutions $y\sim p_{\theta^{\prime}}(y|x)$, achieving the optimal solution (see Fig. 2).
@@ III-B Convergence to the Optimal Solution
In the previous sections, we have successfully transformed the goal into learning and sampling $p(\Theta_{x})$. In this section, we discuss the quality requirements of suboptimal solutions and the generation convergence to optimal solutions.
@@ III-B1 Gap between suboptimal and optimal parameterizations
Let $(x,\theta^{*}),x\in\mathcal{X}$ and $(x,\theta^{\prime}),\theta^{\prime}\sim p(\Theta_{x})$. Although it is not possible to precisely quantify the magnitude of suboptimal perturbations, for the sake of initial discussion, we assume that the perturbations caused by suboptimal solutions are smoothly random and the mean $\mathbb{E}(\Theta_{x})=\mu=\theta^{*}$ holds. With $\mathbb{V}(\Theta_{x})=\sigma^{2}$ as the variance of $p(\Theta_{x})$, we can get Chebyshev’s inequality [49] as below
$p\{|\theta^{\prime}-\theta^{*}|\geq\epsilon\}\leq\frac{\sigma^{2}}{\epsilon^{2}},\ \forall\epsilon>0.$ (9)
Eq. (9) presents the impact of the model solution generation variance and the quality of suboptimal solutions on the gap between the learned suboptimal and optimal parameterizations. The probability that the parametrized $\theta^{\prime}$ of the model output falls outside the $\epsilon$-neighborhood of $\theta^{*}$ can be small. However, since the generation variance of the diffusion model cannot be theoretically quantified, we conduct simulations and analysis in the experimental section to explore this further.
III-B2 Expectation of hitting $y^{*}$
For the $N$-dimensional binary classification problem in the discrete solution space, we have $|\theta_{i}-\theta^{*}_{i}|\!=\!\epsilon,\theta\in\dot{U}(\theta^{*},\epsilon),i\in\{1,...,N\}$. With $\theta^{*}_{i}=1$ generally, the probability of hitting $y^{*}$ in a single sampling is $(1-\epsilon)^{N}$ with independent variables in each dimension. The probability of hitting $y^{*}$ at least once in $n$ samplings can be calculated as $1-[1-(1-\epsilon)^{N}]^{n}$, which is finite.
For the $N$-dimensional regression problem in the continuous solution space, let $\theta\in\dot{U}(\theta^{*},\epsilon)$ have mean $\mu+\epsilon$ and variance $\gamma^{2}$. Keeping the implication of the mean consistent, the parametrized distribution can be independent of the diffusion model sampling. In our discussion, we use Gaussian distribution $\mathcal{N}(\mu+\epsilon,\gamma^{2})$ for feasible solution sampling. Take the tolerable gap as $\delta>0$, the probability of falling $[y^{*}_{i}-\delta,y^{*}_{i}+\delta]$ in a single sampling is
@@ III-B2 Expectation of hitting $y^{*}$
where $y_{i}\sim\mathcal{N}(\mu_{i}+\epsilon,\gamma^{2}_{i}),\mu_{i}=y^{*}_{i}$.
Also with independent variables in each dimension, the probability of hitting $y^{*}$ at least once in $n$ samplings can be calculated as 1($1-(1-\prod^{N}_{i=1}P_{i})^{n}$, which is finite.
Therefore, given an $\epsilon$-neighborhood of $\theta^{*}$ as $\dot{U}(\theta^{*},\epsilon)$, and $\theta$ is taken from $\dot{U}(\theta^{*},\epsilon)$. Let $q(\theta^{n})$ be the probability of hitting $y^{*}$ at least once in $n$ samples $y\sim p_{\theta}(y|x)$, if $\epsilon$ in Eq. (9) makes the gap between $\theta^{\prime}$ and $\theta^{*}$ sufficiently small, a finite number of samples $n$ can make the expectation $\mathbb{E}(q(\theta^{n}))\approx 1$.
@@ III-B3 The lower bound on the number of samples
To demonstrate the impact of $N$ and $n$ on the expectation of hitting $y^{*}$, we provide Fig. LABEL:fig_empirical_ver and give the detailed calculation settings in Sec. V-B. Based on the above discussion, we can give a lower bound on $n$ that makes $\mathbb{E}(q(\theta^{n}))\approx 1$ hold. By simplifying $p_{\epsilon}=p\{|\theta^{\prime}-\theta^{*}|\geq\epsilon\}$ and letting $a>0$ and $a\to 0$, we can multiply both sides of Eq. (9) by $1-[1-(1-\epsilon)^{N}]^{n}$ to obtain
$\begin{split}(1-p_{\epsilon})[1-(1-(1-\epsilon)^{N})^{n}]>\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \\ \ \ \ (1-\frac{\sigma^{2}}{\epsilon^{2}})[1-(1-(1-\epsilon)^{N})^{n}]+a,\ \forall\epsilon>0,\end{split}$ (11)
where $01-[1-(1-\epsilon)^{N}]^{n}>$ 0 does not change the sign of the inequality, and the left side represents the probability that $\theta\in\dot{U}(\theta^{*},\epsilon)$ and $\mathbb{E}(q(\theta^{n}))\approx 1$ occur simultaneously. For a given $a$, $p(\Theta_{x})$ in our problem is almost impossible to be the three-point distribution required by the equality condition in Chebyshev’s inequality111 https://en.wikipedia.org/wiki/Chebyshev%27s_inequality., so the sign $\geq$’ can be directly replaced by ‘$>$’, thus an invalid inequality will not appear because $a\to 0$.
With the Eq. (9) holds, we observe that $p_{\epsilon}\leq\frac{\sigma^{2}}{\epsilon^{2}}$. After transposition and simplification, we can get the lower bound on $n$ for the discrete solution space:
$n>\frac{{\rm ln}(\frac{a}{p_{\epsilon}-\sigma^{2}/\epsilon^{2}}+1)}{{\rm ln}(1-(1-\epsilon)^{N})},\ \forall\epsilon>0.$ (12)
For a continuous solution space, we use 1($1-(1-\prod^{N}_{i=1}P_{i})^{n}$ to replace $1-(1-(1-\epsilon)^{N})^{n}$ and the lower bound can be calculated as
$n>\frac{{\rm ln}(\frac{a}{p_{\epsilon}-\sigma^{2}/\epsilon^{2}}+1)}{{\rm ln}(1-\prod^{N}_{i=1}P_{i})},\ \forall\epsilon>0.$ (13)
To demonstrate the impact of $N$ on the lower bound on $n$ that supports $\mathbb{E}(q(\theta^{n}))\approx 1$, we illustrate it in Fig. LABEL:fig_empirical_ver and the detailed settings in Sec. V-B.
Based on the above discussion, we explain that the model is capable of learning a solution distribution $\theta^{\prime}$ that is sufficiently close to the optimal parametrization $\theta^{*}$. Furthermore, within a finite number of samples, the expectation of hitting the optimal solution at least once can approach $1$. Exploratively, we provide a lower bound expression (Eq. (12), Eq. (13)) for the number of samples $n$ that supports this expectation approaching $1$. These equations are further studied by experiments and model results, thereby providing valuable basis and reference conditions.

### v3

@@ II-B2 Offloading execution cost
$C^{o}_{i}=\alpha_{i}\frac{V_{i}}{F}+(1-\alpha_{i})P_{e}\frac{V_{i}}{F},\ \forall i\in\{1,...,L\}.$ (6)
Therefore, the total offloading cost of $E_{i}$ can be calculated as $D_{i}(C^{t}_{i}+\frac{C^{o}_{i}}{A_{i}})$.
Since the computing task has a maximum tolerable delay $\tau_{i}$ [seconds], let $\rho_{i}\in[0,1]$ represent the minimum computational resource allocation ratio required for offloading execution without timeout, and let $\psi_{i}\in\{0,1\}$ represent whether the local execution will timeout or not. Only if the offloading was able to process without a timeout, $\rho_{i}>0$. As a result, $\rho_{i}=0$ if $\frac{I_{i}}{r^{u}_{i}}+\frac{V_{i}}{F}>\tau_{i}$ and $\rho_{i}=\frac{V_{i}}{(\tau_{i}-I_{i}/r^{u}_{i})F}$ if $\frac{I_{i}}{r^{u}_{i}}+\frac{V_{i}}{F}\leq\tau_{i}$. Furthermore, $\psi_{i}=0$ if the local execution will timeout, $\psi_{i}=1$ otherwise. Therefore, we use $(C^{l}_{i},C^{t}_{i},C^{o}_{i},\rho_{i},\psi_{i})$ as the 5-dimensional edge feature in $\mathcal{G}$ to represent computing tasks, communication parameters, and resource settings.
@@ II-C Weighted Cost Minimization Problem
According to Eqs. (4), (5) and (6), the weighted cost minimization problem can be formulated as
$\mathbb{P}:\underset{\{\mathcal{D},\mathcal{A}\}}{{\rm min}}\ \sum^{L}_{i=1}(1-D_{i})C^{l}_{i}+D_{i}(C^{t}_{i}+\frac{C^{o}_{i}}{A_{i}}),$
@@
Let $\{\mathcal{D}^{*},\mathcal{A}^{*}\}$ be the optimal solution for the given input parameters, the problem $\mathbb{P}$ can be transformed from the original mix-integer non-linear programming (MINLP) to a multi-task graph optimization. It is important to note that although $\mathbb{P}$ is formally an MINLP problem and the existing solvers cannot directly handle such a challenging graph-structured problem. For example, Gurobi and GEKKO solver only accept vector-based problem and constraint formulations, and the time cost for solving MINLP problems is excessively high. In this context, with the node features that represent node types (e.g., $1$ represents a server node and 00 represents a user node) and edge features that explicitly denote objective and constraint, the graph-based formulation can provide rich information and can be transferable to alternative problems.
@@ III Problem Analysis
In this section, we present the principles and concepts of solution distribution learning, explicitly highlighting the robustness of this learning objective with respect to solution quality, thereby solving the optimization problem in MEC networks from the suboptimal dataset.
@@ III-A Why and How to Learn the Solution Distribution?
For a network optimization problem, existing methods [38, 25, 24, 26, 37, 5, 13] generally obtain truth input-solution paired samples such as $(\mathbf{x},\mathbf{y}^{*}),\mathbf{x}\in\mathcal{X},\mathbf{y}^{*}\in\mathcal{Y}_{\mathbf{x}}$, $\mathcal{X}$ and $\mathcal{Y}_{\mathbf{x}}$ represent the input parameter space and the feasible solution space for a given $\mathbf{x}$, respectively. Inspired by [58], we extend the definition of parametrized solution space and parametrize the solution space with a continuous vector $\boldsymbol{\theta}\in\mathbb{R}^{N}$, where $N$ represents the dimension of the optimization variable. This probability of each feasible solution $\mathbf{y}$ can be estimated as
$p_{\boldsymbol{\theta}}(\mathbf{y}|\mathbf{x})\propto{\rm exp}(\sum^{N}_{i=1}y_{i}\theta_{i})\ \ s.t.\ \mathbf{y}\in\mathcal{Y}_{\mathbf{x}},\ \mathbf{x}\in\mathcal{X},$ (8)
@@
where $p_{\boldsymbol{\theta}}$ is an energy function indicating the probability of each solution over the feasible solution space. We first consider the discrete solution space that $\mathbf{y}$ is an $N$-dimensional vector with $y_{i}\in\{0,1\}$, and the higher value of $\theta_{i}$ denotes a higher probability that $y_{i}$ is consistent with $\mathbf{y}^{*}$. There exists $\boldsymbol{\theta}^{*}$ that is the optimal parametrization of the solution space for a given $\mathbf{x}\in\mathcal{X}$ if $\boldsymbol{\theta}^{*}$ satisfies $p_{\boldsymbol{\theta}^{*}}(\mathbf{y}^{*}|\mathbf{x})=1$.
@@
Accordingly, $\boldsymbol{\theta}$ represents a solution distribution. Taking the discrete binary classification problem as an example, for the optimal solution $\mathbf{y}^{*}$, it is parameterized using a binomial distribution. Specifically, $\theta^{*}_{i0}=0$ and $\theta^{*}_{i1}=1$ if $y^{*}_{i}=1$, while $\theta^{*}_{i0}=1$ and $\theta^{*}_{i1}=0$ if $y^{*}_{i}=0$. Here, $\theta^{*}_{i0}$ represents the probability that $y_{i}=0$, and $\theta^{*}_{i1}$ represents the probability that $y_{i}=1$. Therefore, $(\mathbf{x},\mathbf{y}^{*})$ can be replaced by $(\mathbf{x},\boldsymbol{\theta}^{*})$ for the purpose of solution learning.
@@
Although learning $\boldsymbol{\theta}^{*}$ can guarantee accurate results, it is clear that $\boldsymbol{\theta}^{*}$ is overly specific, making it indistinguishable from discriminative training and causing the parameterization to be redundant. For the NP-hard problems addressed in this paper, efficiently obtaining a large amount of ground-truth samples for discriminative supervised training is not feasible. In contrast, suboptimal solutions are significantly easier to obtain, which makes the solution distribution learning with the generative diffusion model possible.
@@
Let $(\mathbf{x},\mathbf{y}^{\prime})$ be a suboptimal solution for a given $\mathbf{x}$, and $(\mathbf{x},\boldsymbol{\theta}^{\prime})$ be the parametrization. To avoid verbosity, we will not introduce the definition of the gap between suboptimal solutions and the optimal solution here and leave it in Sec. III-B. We know that suboptimal parameterizations are not unique [^2] For example, taking $\boldsymbol{\theta}^{\prime}_{1}=(0.9,0.1),\boldsymbol{\theta}^{\prime}_{2}=(0.95,0.05)$ for $\boldsymbol{\theta}^{*}=(1,0)$, where repeated sampling of $p_{\boldsymbol{\theta}^{\prime}_{1}}(\mathbf{y}|\mathbf{x})$ or $p_{\boldsymbol{\theta}^{\prime}_{2}}(\mathbf{y}|\mathbf{x})$ can easily hit an optimal solution.. Therefore, let $Y^{\prime}_{\mathbf{x}}$ be the set of suboptimal solutions for $\mathbf{x}$, and $\boldsymbol{\Theta}_{\mathbf{x}}$ be the set of suboptimal parameterizations which contains $\boldsymbol{\theta}^{*}$.
@@
For the continuous solution space, it is impossible to make the calculated solution completely equal to $\mathbf{y}^{*}$ because of the infinite precision. As a discussion example, we parametrize $\mathbf{y}^{*}$ using sharp Laplace distribution, making the probability of the optimal solution approximate to 1. For example, for $y^{*}_{i}$, $\theta^{*}_{i}=y^{*}_{i}$ for ${\rm Laplace}(\theta^{*}_{i},{\rm scale\_factor})$, smaller ${\rm scale\_factor}$ means sharper probability density function and the probability of $\mathbf{y}^{*}$ is the largest and tends to 1. Thus, we can obtain the continuous solution space’s $Y^{\prime}_{\mathbf{x}}$ and $\boldsymbol{\Theta}_{\mathbf{x}}$.
@@
Consequently, when given only a dataset of suboptimal solutions, learning the parametrization of the solution space with parallel sampling can lead to the optimal solution. Specifically, the model aims to learn $p(\boldsymbol{\Theta}_{\mathbf{x}})$ to obtain $\boldsymbol{\theta}^{\prime}\!\sim\!p(\boldsymbol{\Theta}_{\mathbf{x}})$ and sample feasible solutions $\mathbf{y}\sim p_{\boldsymbol{\theta}^{\prime}}(\mathbf{y}|\mathbf{x})$ to achieve the optimal solution, as shown in Fig. 2.
@@ III-B Convergence to the Optimal Solution
In the previous section, we successfully transformed the goal into learning and sampling $p(\boldsymbol{\Theta}_{\mathbf{x}})$. In this section, we discuss the quality requirements of suboptimal solutions and the generation convergence to optimal solutions.
@@ III-B1 Gap between suboptimal and optimal parameterizations
Let $(\mathbf{x},\boldsymbol{\theta}^{*}),\mathbf{x}\in\mathcal{X}$ and $(\mathbf{x},\boldsymbol{\theta}^{\prime}),\boldsymbol{\theta}^{\prime}\sim p(\boldsymbol{\Theta}_{\mathbf{x}})$. Although it is not possible to precisely quantify the magnitude of suboptimal perturbations, for the sake of initial discussion, we assume that the perturbations caused by suboptimal solutions are smoothly random and the mean $\mathbb{E}(\boldsymbol{\Theta}_{\mathbf{x}})=\boldsymbol{\mu}=\boldsymbol{\theta}^{*}$ holds. With $\mathbb{V}(\boldsymbol{\Theta}_{\mathbf{x}})=\boldsymbol{\sigma}^{2}$ as the variance of $p(\boldsymbol{\Theta}_{\mathbf{x}})$, we can obtain Chebyshev’s inequality [59] as below
$p\{|\boldsymbol{\theta}^{\prime}-\boldsymbol{\theta}^{*}|\geq\boldsymbol{\epsilon}\}\leq\frac{\boldsymbol{\sigma}^{2}}{\boldsymbol{\epsilon}^{2}},\ \forall\boldsymbol{\epsilon}>0.$ (9)
Eq. (9) presents the impact of the model solution generation variance and the quality of suboptimal solutions on the gap between the learned suboptimal and optimal parameterizations. By controlling the variance of generation in Eq. (9), the probability that the parametrized $\boldsymbol{\theta}^{\prime}$ of the model output falls outside the $\boldsymbol{\epsilon}$-neighborhood of $\boldsymbol{\theta}^{*}$ can be small. However, since the generation variance of the diffusion model cannot be theoretically quantified, we conduct simulations and analysis in the experimental section to explore this further.
III-B2 Expectation of hitting $\mathbf{y}^{*}$
For the $N$-dimensional binary classification problem in the discrete solution space, we have $|\theta_{i}-\theta^{*}_{i}|\!=\!\boldsymbol{\epsilon},\boldsymbol{\theta}\in\dot{U}(\boldsymbol{\theta}^{*},\boldsymbol{\epsilon}),i\in\{1,...,N\}$, where $\dot{U}(\boldsymbol{\theta}^{*},\boldsymbol{\epsilon})$ represents the center $\boldsymbol{\theta}^{*}$ and the neighborhood with radius $\boldsymbol{\epsilon}$. With $\theta^{*}_{i}=1$ generally, the probability of hitting $\mathbf{y}^{*}$ in a single sampling is $(1-\boldsymbol{\epsilon})^{N}$ with independent variables in each dimension. The probability of hitting $\mathbf{y}^{*}$ at least once in $n$ samplings can be calculated as $1-[1-(1-\boldsymbol{\epsilon})^{N}]^{n}$, which is finite.
For the $N$-dimensional regression problem in the continuous solution space, let $\boldsymbol{\theta}\in\dot{U}(\boldsymbol{\theta}^{*},\boldsymbol{\epsilon})$ have mean $\boldsymbol{\mu}+\boldsymbol{\epsilon}$ and variance $\boldsymbol{\gamma}^{2}$. Keeping the implication of the mean consistent, the parametrized distribution can be independent of the diffusion model sampling. In our discussion, we use Gaussian distribution $\mathcal{N}(\boldsymbol{\mu}+\boldsymbol{\epsilon},\boldsymbol{\gamma}^{2})$ for feasible solution sampling. Take the tolerable gap as $\delta>0$, the probability of falling $[y^{*}_{i}-\delta,y^{*}_{i}+\delta]$ in a single sampling is
@@ III-B2 Expectation of hitting $\mathbf{y}^{*}$
where $y_{i}\sim\mathcal{N}(\mu_{i}+\boldsymbol{\epsilon},\gamma^{2}_{i}),\mu_{i}=y^{*}_{i}$.
Also with independent variables in each dimension, the probability of hitting $\mathbf{y}^{*}$ at least once in $n$ samplings can be calculated as 1($1-(1-\prod^{N}_{i=1}P_{i})^{n}$, which is finite.
Given an $\boldsymbol{\epsilon}$-neighborhood of $\boldsymbol{\theta}^{*}$ as $\dot{U}(\boldsymbol{\theta}^{*},\boldsymbol{\epsilon})$, and $\boldsymbol{\theta}$ is taken from $\dot{U}(\boldsymbol{\theta}^{*},\boldsymbol{\epsilon})$. Let $q(\boldsymbol{\theta}^{n})$ be the probability of hitting $\mathbf{y}^{*}$ at least once in $n$ samples $\mathbf{y}\sim p_{\boldsymbol{\theta}}(\mathbf{y}|\mathbf{x})$, if $\boldsymbol{\epsilon}$ in Eq. (9) makes the gap between $\boldsymbol{\theta}^{\prime}$ and $\boldsymbol{\theta}^{*}$ sufficiently small, a finite number of solution sampling $n$ can make the expectation $\mathbb{E}(q(\boldsymbol{\theta}^{n}))\approx 1$.
@@ III-B3 The lower bound on the number of samples
To demonstrate the impact of $N$ and $n$ on the expectation of hitting $\mathbf{y}^{*}$, we provide Fig. LABEL:fig_empirical_ver and give the detailed calculation settings in Sec. V-B. Based on the above discussion, we can give a lower bound on $n$ that makes $\mathbb{E}(q(\boldsymbol{\theta}^{n}))\approx 1$ hold. By simplifying $p_{\boldsymbol{\epsilon}}=p\{|\boldsymbol{\theta}^{\prime}-\boldsymbol{\theta}^{*}|\geq\boldsymbol{\epsilon}\}$ and letting $\mathbf{a}>\mathbf{0}$ and $\mathbf{a}\to\mathbf{0}$, we can multiply both sides of Eq. (9) by $1-[1-(1-\boldsymbol{\epsilon})^{N}]^{n}$ to obtain
$\begin{split}(1-p_{\boldsymbol{\epsilon}})[1-(1-(1-\boldsymbol{\epsilon})^{N})^{n}]>\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \\ \ \ \ (1-\frac{\boldsymbol{\sigma}^{2}}{\boldsymbol{\epsilon}^{2}})[1-(1-(1-\boldsymbol{\epsilon})^{N})^{n}]+\mathbf{a},\ \forall\boldsymbol{\epsilon}>0,\end{split}$ (11)
where $01-[1-(1-\boldsymbol{\epsilon})^{N}]^{n}>$ 0 does not change the sign of the inequality, and the left side represents the probability that $\boldsymbol{\theta}\in\dot{U}(\boldsymbol{\theta}^{*},\boldsymbol{\epsilon})$ and $\mathbb{E}(q(\boldsymbol{\theta}^{n}))\approx 1$ occur simultaneously. For a given $\mathbf{a}$, $p(\boldsymbol{\Theta}_{\mathbf{x}})$ in our problem is almost impossible to be the three-point distribution required by the equality condition in Chebyshev’s inequality [59], so the sign $\geq$’ can be directly replaced by ‘$>$’, thus an invalid inequality will not appear because $\mathbf{a}\to\mathbf{0}$.
With the Eq. (9) holds, we observe that $p_{\boldsymbol{\epsilon}}\leq\frac{\boldsymbol{\sigma}^{2}}{\boldsymbol{\epsilon}^{2}}$. After transposition and simplification, we can get the lower bound on $n$ for the discrete solution space:
$n>\frac{{\rm ln}(\frac{\mathbf{a}}{p_{\boldsymbol{\epsilon}}-\boldsymbol{\sigma}^{2}/\boldsymbol{\epsilon}^{2}}+1)}{{\rm ln}(1-(1-\boldsymbol{\epsilon})^{N})},\ \forall\boldsymbol{\epsilon}>0.$ (12)
For a continuous solution space, we use 1($1-(1-\prod^{N}_{i=1}P_{i})^{n}$ to replace $1-(1-(1-\boldsymbol{\epsilon})^{N})^{n}$ and the lower bound can be calculated as
$n>\frac{{\rm ln}(\frac{\mathbf{a}}{p_{\boldsymbol{\epsilon}}-\boldsymbol{\sigma}^{2}/\boldsymbol{\epsilon}^{2}}+1)}{{\rm ln}(1-\prod^{N}_{i=1}P_{i})},\ \forall\boldsymbol{\epsilon}>0.$ (13)
To make $\mathbb{E}(q(\boldsymbol{\theta}^{n}))\approx 1$ hold, the lower bound of the number of solution sampling $n$ is given by (12) and (13), respectively. Therefore, the diffusion model with solution sampling $n$ is capable of learning a solution distribution $\boldsymbol{\theta}^{\prime}$, which is sufficiently close to the optimal parametrization $\boldsymbol{\theta}^{*}$. To demonstrate the impact of $N$ on the lower bound on $n$ that supports $\mathbb{E}(q(\boldsymbol{\theta}^{n}))\approx 1$, we illustrate it in Fig. LABEL:fig_empirical_ver and the detailed settings in Sec. V-B.